# Fewer Hallucinations, More Verification: A Three-Stage LLM-Based Framework for ASR Error Correction

### v1

@@ I Introduction
Text error correction has become an important component in modern speech recognition systems to reduce error rates [1, 2, 3, 4]. Early approaches rely on shallow neural networks or rule-based strategies to handle simple error patterns [5, 6].
Recent advances in deep learning have given rise to two dominant paradigms: autoregressive sequence-to-sequence methods [7, 8, 9] and non-autoregressive edit-based methods [1, 2]. Autoregressive sequence-to-sequence methods that take advantage of encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [10], include translation-style correction frameworks [7, 8] and entity-aware transformers [9]. [8] treats error correction task as translation task, while [7] additionally takes into account domain adaptation; non-autoregressive edit-based methods like FastCorrect series [1, 2] that predict edit operations via duration modeling. FastCorrect [1] first introduced the concept of edit distance predictors and correctors, while FastCorrect2 [2] also utilizes multiple candidate sentences generated by ASR for correction. SoftCorrect [4] combines both approaches, providing clear error detection signals while utilizing CTC loss and sequence-to-sequence models.
While conventional models require collecting large amounts of labeled data and training for specific tasks, the rise of LLMs offers new opportunities through their contextual adaptability to perform various language-related subtasks including text classification [11], information extraction [12], dialogue systems [13], and so on.
The application of LLMs in NLP subtasks typically follows two paradigms: prompt-based methods and fine-tuning methods. Research on prompt-based methods focuses on how to stimulate the potential capabilities of LLMs , such as using generative prompts for common-sense reasoning [14], and enhancing reasoning abilities through Chain of thought Prompts [15]. Research on model fine-tuning usually employs LoRA [16] technology to improve performance on specific tasks. But it still a little resource consuming. Directly using general large language models for correction, instead of fine-tuning task-specific models, provides a more convenient approach.
Building on this line of research, prior studies have explored incorporating LLMs during the correction process[17, 18, 19], achieving notable improvements and reductions in WER. Unfortunately, the above methods either provide additional information or fine-tune LLM in a specific domain, which limits the error correction in a wide range of real-world domains.
A key question arises: Is it possible to use a general LLM to correct ASR recognition errors directly without providing additional information or collecting a large number of error-correction text pairs to finetune the model?
However, directly using general LLM without making extra efforts to correct speech recognition errors will face many hallucination problems.[20].
Figure 1: Reliable LLM Correction Framework
We further confirm this with experiments and propose the RLLM-CF, a novel LLM-based framework to address these challenges.
Compared with other existing LLM-based methods, we do not use other information or fine-tune the LLM, but directly use the original knowledge of the general LLM, which makes a general correction model can be used in any field without any effort. And, the bigger difference is that we do not simply adopt the modified results of LLM directly, but use a series of procedures to combat the illusion of LLM. In order to reject the hallucination problem of LLM, we adopt error pre-detection and reasonning process verification. At the same time in order to limit the output space of LLM, we provide a CoT that breaks down the error correction process into multiple subtasks, so that the correction process of LLM is logical and each correction has a basis.
The effectiveness of our approach in boosting recognition accuracy is clearly demonstrated through experiments on the AISHELL-1, AISHELL-2, and Librispeech datasets. Specifically, the best-performing GPT-4o model, when enhanced with our method, achieves a remarkable reduction in the original error rates: 21% on AISHELL-1, 11% on AISHELL-2, and 9% and 11.4% on the clean and other subsets of Librispeech, respectively.
@@
The remainder of this paper is organized as follows: Section II reviews recent approaches to ASR correction based on LLMs, with a particular focus on the challenges posed by hallucination issues inherent to LLMs. Section III provides a detailed description of our proposed framework. Section IV presents comparative experiments on benchmark datasets, followed by a framework analysis in Section V. Section VI summarizes the paper.
@@ II-A LLM-based ASR Error Correction
In recent years, the integration of LLMs into ASR error correction pipelines has garnered increasing attention. Min and Wang [20] explores the direct application of LLM to error correction, and concluded that it is not feasible to directly correct using models like GPT-4o due to hallucinations problem. Ma et al. [19] explored the results of combining N-best Rescore and LLM correction. However, it is difficult to obtain n-best infomation after recognizing speeches. Yang et al. [18] compares the difference between using LLM rescore method and the generated method. After providing domain-specific information or fine-tuning the LLM, the generated method outperforms the rescore method. However, it requires specific domain information or fine-tuning the LLM. In contrast to existing approaches, our method eliminates the need for supplementary resources (e.g., n-best lists or domain-specific knowledge), requires no LLM fine-tuning, and maintains critical verification of the model’s correction outputs without blind reliance.
@@ II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to situations in which the generated response is grammatically correct, fluent, and plausible, but deviates from the input of the source or contradicts the factual accuracy[21, 22]. When LLMs are used for ASR error correction, the errors produced by the model are closely related to hallucinations. Min and Wang [20] conducted an empirical study on integrating LLM directly into ASR systems to improve transcription quality. Their experiments confirmed that the integration of powerful generative models can inadvertently introduce hallucination errors, where the model fabricates content not present in the original audio.
In this work, we further confirm and analyze the hallucination phenomenon that arises when LLMs are directly applied for ASR error correction with simple prompts in IV-A. Moreover, our proposed method effectively mitigates this phenomenon, leading to more reliable and accurate error correction.
@@ III Reliable ASR Error Correction Framework with LLM
In this section, we first present the overall architecture of the RLLM-CF in III-A. After that, we mention the error pre-detection in III-B. Next, we show how to decompose speech recognition error correction into sub-tasks through chain of thought in III-C. Finally, we introduce the answer verification in III-D.
@@ III-A Reliable LLM Correction Framework
The Reliable LLM Correction Framework is showed in Figure 1. The blue part is executed by the LLM, and the yellow part is executed by the program. The core part of the framework is mainly composed of three components: Error Pre-Detection, Chain of Thought Sub-Tasks Iterative Correction, and Reason Process Verification. The detailed prompt design is shown in the Figure III-D. The text in the white boxes represents the specific prompt content, while the blue boxes provide explanations. It is important to note that we also provide 3 additional examples for correction, including the specific reasoning process. Due to space constraints, we did not show them in the figure. The entire algorithm is described in Alg 1.
Figure 2: Prompt Design
@@ III-B Error Pre-Detection
Regarding the hallucination issues demonstrated in Table I, general LLMs typically generate more errors through hallucinations than they correct. This necessitates establishing ”error prevention priority” as the fundamental principle for correction tasks. To prevent LLMs from modifying correct text into incorrect text, we first instruct the LLM to detect errors in the input sentence. If no errors are found, it directly outputs the sentence; otherwise, it proceeds to the error correction stage.
III-C Chain of Thought Sub-Tasks Iterative Correction
In the error correction stage, the LLMs often fail to infer the correct answer or generate hallucinations, making arbitrary modifications to the recognized sentences without a reliable basis. Therefore, we break down the speech recognition error correction task into four sub-tasks, localization, pronunciation, candidate generation, and selection as shown in Figure 3—using a chain-of-thought approach to enhance the reliability of the LLMs’ reasoning. Even answers generated through the chain-of-thought reasoning may not be fully accurate and could still contain multiple errors. Thus, we conduct a confidence assessment of the answer generated by the LLMs. If the LLM maintains confidence in the answer, it directly outputs the corrected answer in the required format; otherwise, it continues to iterate through the error correction process as described above before the maximum number of allowed corrections is exceeded.
Figure 3: Sub-Tasks Example
@@ III-D Answer Verification
After getting the final answer, we add verification steps to ensure the LLM followed instructions correctly. Specifically, we check two things using the LLM’s self-correction feedback: 1) whether the answer matches the required format, and 2) whether it completed all necessary reasoning steps. Only when both checks pass do we use the corrected result; otherwise, we keep the original answer.
@@ IV Experiment
The organization of the experimental section of this study is as follows: After a comprehensive study and comfirmation on LLM Hallucination in IV-A, we present a comprehensive introduction to the experimental setup in Section IV-B and Section IV-C. Subsequently, we exhibit the experimental results of the models on the AISHELL-1 [23], AISHELL-2 [24], and Librispeech datasets [25] in Section IV-D. Finally, we conduct an overall analysis of the results and perform an analysis of the noun recall rate in IV-E.
@@ IV-A LLM Hallucination
Based on our experiments with plain prompt, we can roughly divide LLM Hallucinations in ASR error correction into the following types: faithful hallucinations ( instruction violation, redundant output, continue writing, blank output and grammar correction ) and factual hallucinations ( make mistake ) as shown in Tabel I.
IV-B Dataset Selection
TABLE II: AISHELL-1 Results
Model Decode LLM Correct Error Metrics Change CER Noun Recall
sub del insert
Without LLM Correction
conformer attention - 4.65 0.29 0.12 - 5.06 79.32
conformer attention rescore - 4.42 0.12 0.08 - 4.62 79.93
conformer ctc greedy search - 4.95 0.13 0.09 - 5.17 77.91
conformer ctc prefix search - 4.95 0.13 0.09 - 5.17 77.88
With GPT-4o Correction
conformer attention GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
conformer attention rescore GPT-4o 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
conformer ctc greedy search GPT-4o 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
conformer ctc prefix search GPT-4o 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
With DeepSeek v2 Correction
conformer attention rescore DeepSeek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
conformer attention rescore DeepSeek v2 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
conformer ctc greedy search DeepSeek v2 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
conformer ctc prefix search DeepSeek v2 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
TABLE III: AISHELL-2 Results
Model Decode Chunk Size LLM Correct Error Metrics Change CER Noun Recall
sub del insert
conformer attention rescore 16 - 5.26 0.17 0.14 - 5.57 77.64
conformer attention rescore 16 GPT-4o 4.47 0.29 0.19 -0.62(11%) 4.95 79.66
TABLE IV: Librispeech Results (test-clean/other)
Model Decode LLM Correct Error Metrics Change WER Noun Recall
sub del insert
Without LLM Correction
conformer attention - 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
conformer attention rescore - 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
conformer ctc greedy search - 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
conformer ctc prefix search - 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
With GPT-4o Correction
conformer attention GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
conformer attention rescore GPT-4o 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
conformer ctc greedy search GPT-4o 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
conformer ctc prefix search GPT-4o 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
@@ IV-B Dataset Selection
We select three datasets as our evaluation standards, including the Chinese datasets AISHELL-1[23] and AISHELL-2 [24], and the English dataset LibriSpeech [25], aimming to conducting experimental assessments on both Chinese and English data. LibriSpeech test set is divided into two subsets: test-clean with high quality and test-other with noisy audio.
IV-C Model Setting
@@
We train a Conformer baseline model using the Wenet tool [26] based on the U2++ method across three datasets. For the AISHELL-1 and Librispeech datasets, we employ a non-streaming model architecture and test four different decoding strategies: Attention, Attention Rescore, CTC Greedy Search, and CTC Prefix Search. To further validate the performance of streaming models, we have conducted experiments on the AISHELL-2 dataset using the Attention Rescore decoding method with a chunk size of 16.
The Conformer model is trained using the Wenet tool [26] and trained separately on different datasets. The encoder is designed with a Conformer structure consisting of 12 blocks, each equipped with an attention dimension of $256$, $4$ attention heads, and $2048$ linear units. It incorporates relative position encoding and utilizes the Swish activation function. The decoder is composed of a bitransformer with three forward and three backward blocks. The network is a hybrid CTC/attention model, with the CTC weight adjusted to $0.3$.
@@
As for LLMs, we have chosen GPT-4o [27] and Deepseek v2 [28] , all of which have strong capabilities in handling both Chinese and English. The temperature parameter was set to $0.2$, and the top_p parameter to $0.8$.
@@
TABLE V: Framework Analysis Results on AISHELL-1 with DeepSeek v2 and Attention Decoding
LLM Correct Error Metrics CER Change Noun Recall Cost Tokens (in/out)
sub del insert

### v2

@@ I Introduction
Text error correction has become an essential component in modern speech recognition systems, aiming to reduce word error rates (WER) [1, 2, 3, 4]. Early methods typically relied on shallow neural networks or rule-based approaches to handle simple error patterns [5, 6]. Recent advances in deep learning have led to two dominant paradigms: autoregressive sequence-to-sequence models [7, 8, 9] and non-autoregressive edit-based models [1, 2]. Autoregressive models exploit encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [10], including translation-style correction frameworks [7, 8] and entity-aware transformers [9]. In parallel, non-autoregressive edit-based models such as FastCorrect [1, 2] and SoftCorrect [4] predict edit operations through duration modeling and integrate CTC loss with sequence-to-sequence frameworks to enhance error detection.
While conventional approaches depend heavily on large labeled datasets and task-specific training, the emergence of large language models (LLMs) offers new opportunities due to their strong contextual adaptability. LLMs have demonstrated remarkable performance across a wide range of language-related tasks, including text classification [11], information extraction [12], and dialogue systems [13]. This raises a fundamental question: Can a general-purpose LLM correct ASR errors directly without fine-tuning or additional information?
Current LLM applications typically follow two paradigms: prompt-based methods and fine-tuning strategies. Prompt-based methods aim to stimulate latent capabilities, employing techniques such as generative prompting for commonsense reasoning [14] and Chain-of-Thought (CoT) prompting to enhance reasoning [15]. Fine-tuning approaches, often using low-rank adaptation (LoRA) [16], improve task-specific performance but remain resource-intensive. Building on the aforementioned studies, recent works have explored LLM integration into ASR error correction pipelines [17, 18, 19], achieving notable CER/WER reductions. However, these methods typically require either domain-specific fine-tuning or additional contextual information, which constrains their scalability in real-world applications. Therefore, leveraging the general knowledge embedded in LLMs without fine-tuning or external inputs emerges as a more practical and scalable alternative. Nevertheless, directly applying general LLMs often results in hallucination issues [20], posing a significant challenge to reliable correction.
Figure 1: Overview of the Reliable LLM Correction Framework (RLLM-CF). The blue components represent modules executed by the LLM, while the yellow components are handled by the external program. The green block indicates the recognition result, and the black block denotes the final output. The entire error correction process follows the direction of the arrows.
To address these issues, we propose Reliable LLM Correction Framework (RLLM-CF), a novel LLM-based framework that directly leverages the pretrained knowledge of general LLMs without fine-tuning or external information. RLLM-CF mitigates hallucinations by integrating error pre-detection, reasoning-based verification, and a CoT strategy that structures correction into explainable and grounded subtasks, enabling broad applicability across domains. The proposed framework is validated on AISHELL-1, AISHELL-2, LibriSpeech test-clean and test-other datasets, achieving CER/WER reductions of 21%, 11%, 9%, and 11.4%, respectively.
The main contributions of this paper are as follows:
  • We propose RLLM-CF, an ASR error correction framework base on general-purpose LLM that requires neither fine-tuning nor external information.
  • We introduce a hallucination mitigation mechanism based on error pre-detection and reasoning-based verification, addressing the limitations of prior methods that directly accept LLM correction outputs and risk hallucination errors.
  • We employ a Chain-of-Thought (CoT) prompting strategy to enhance correction reliability by decomposing the correction process into subtasks: localization, pronunciation assessment, candidate generation, and candidate selection.
@@
The remainder of this paper is organized as follows. Section II reviews recent LLM-based ASR correction approaches, with a particular focus on hallucination challenges. Section III details the proposed framework. Section IV presents experimental results and analysis on benchmark datasets. Section V provides further discussions, and Section VI concludes the paper.
@@ II-A LLM-based ASR Error Correction
In recent years, integrating LLMs into ASR error correction pipelines has attracted increasing attention. Min and Wang [20] investigated the direct application of LLMs for error correction and concluded that models such as GPT-4o are ineffective due to hallucination issues. Ma et al. [19] explored combining N-best rescoring with LLM-based correction; however, obtaining the N-best list from the ASR system may impose additional costs in practice. Yang et al. [18] compared LLM-based rescoring and generation methods, demonstrating that the latter outperforms the former when domain-specific information is provided or the LLM is fine-tuned. Nevertheless, such approaches require either domain knowledge or fine-tuning.
In contrast to existing methods, our approach eliminates the need for supplementary resources (e.g., N-best lists or domain-specific knowledge), avoids fine-tuning the LLM, and incorporates critical verification mechanisms to ensure correction reliability without blind reliance on model outputs.
@@ II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to instances where the generated responses, while grammatically correct, fluent, and plausible, deviate from the input or contradict factual information [21, 22]. In the context of ASR error correction, model-induced hallucinations are closely linked to transcription errors. Min and Wang [20] conducted an empirical study on directly integrating LLMs into ASR systems, confirming that powerful generative models can inadvertently introduce hallucination errors by fabricating content absent from the original audio.
In this work, we further validate and analyze the hallucination phenomenon arising from direct LLM-based ASR error correction using simple prompts, as detailed in Section IV-A. Moreover, the proposed method effectively mitigates hallucinations, resulting in more reliable and accurate correction performance.
@@ III Reliable ASR Error Correction Framework with LLM
This section presents the overall architecture of RLLM-CF, including the design of the error pre-detection module, the decomposition of ASR error correction into subtasks via a Chain-of-Thought strategy, and the implementation of an answer verification mechanism. Detailed descriptions of each component are provided in Sections III-A, III-C and III-D.
@@ III-A Reliable LLM Correction Framework
The overall framework of RLLM-CF is illustrated in Figure 1. It consists of three main components: Error Pre-Detection, Chain-of-Thought Subtask Iterative Correction, and Reasoning Process Verification. The prompt design, illustrated in Figure III-D, incorporates three additional few-shot correction examples demonstrating specific reasoning steps, which are omitted for clarity. The complete correction algorithm is outlined in Algorithm 1.
Figure 2: The above picture introduces the specific parts of Prompt’s design in detail.;The text in the white boxes represents the specific prompt content, while the blue boxes provide explanations.
@@ III-B Error Pre-Detection
As shown in Table I, general LLMs often generate hallucinations, producing more errors than corrections. This highlights the necessity of adopting an error prevention first strategy for correction tasks. To prevent LLMs from altering correct content, we first instruct the model to detect errors in the input sentence. If no errors are detected, the sentence is directly retained; otherwise, the model proceeds to the correction stage, referred to as Stage 1 in Algorithm 1.
III-C Chain of Thought Subtask Iterative Correction
During error correction, denoted as Stage 2 in Algorithm 1, LLMs may fail to infer the correct answer or introduce hallucinations by making unwarranted modifications. To address this, we decompose the correction task into four subtasks—localization, pronunciation assessment, candidate generation, and candidate selection—following a CoT strategy to improve reasoning reliability, as illustrated in Figure 3.
Despite this structured reasoning, the generated results may still contain errors. To mitigate this, we perform a confidence evaluation on the generated answer. If the model remains confident, the corrected answer is output in the required format; otherwise, the correction process iterates until either a valid answer is obtained or the maximum number of iterations is reached.
@@ III-D Answer Verification
Following the correction, a verification step is conducted to ensure compliance with task instructions. Specifically, we employ the model’s self-verification feedback to assess: (1) whether the answer conforms to the required format, and (2) whether all reasoning steps are correctly completed. Only when both criteria are satisfied is the corrected output accepted; otherwise, the original input is retained.
Figure 3: The above is a detailed answer to illustrate the composition and specific answers of the subtasks.
@@ IV Experiment
The experimental section is organized as follows. Section IV-A investigates and validates LLM hallucination phenomena. Section IV-B and Section IV-C describe the experimental setup, including dataset selection and model configurations. Experimental results on AISHELL-1 [23], AISHELL-2 [24], and LibriSpeech [25] are reported in Section IV-D. Finally, Section IV-E presents an overall performance analysis, including a detailed evaluation of noun recall rates.
@@ IV-A LLM Hallucination
In experiments with basic prompting, we categorize LLM hallucinations in ASR error correction into two types: faithful hallucinations—including instruction violations, redundant outputs, continuation of writing, blank outputs, and grammar corrections—and factual hallucinations, characterized by content errors. A detailed categorization is summarized in Table I.
IV-B The Dataset Configuration
TABLE II: Character Error Rate (CER) and Noun Recall of GPT-4o and Deepseek v2 on AISHELL-1 ASR Error Correction;The ASR model is a conformer-based AED model trained on AISHELL-1.
Decode Method LLM Correction Error Metrics Performance Metrics
sub del insert Change CER Noun Recall
Attention[26] 4.65 0.29 0.12 - 5.06 79.32
Attention Rescore[27] 4.42 0.12 0.08 - 4.62 79.93
CTC Greedy Search[10] 4.95 0.13 0.09 - 5.17 77.91
CTC Prefix Search[10] 4.95 0.13 0.09 - 5.17 77.88
Attention[26] GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
Attention Rescore[27] 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
CTC Greedy Search [10] 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
CTC Prefix Search[10] 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
Attention[26] Deepseek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
Attention Rescore[27] 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
CTC Greedy Search[10] 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
CTC Prefix Search[10] 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
TABLE III: Word Error Rate (WER) and Noun Recall of GPT-4o on Librispeech ASR Error Correction (test-clean/other Subsets).The ASR Model is conformer-based AED trained on Librispeech.
Decode Method LLM Correction Error Metrics Performance Metrics
sub del insert Change WER Noun Recall
Attention[26] 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
Attention Rescore[27] 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
CTC Greedy Search[10] 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
CTC Prefix Search[10] 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
Attention[26] GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
Attention Rescore[27] 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
CTC Greedy Search[10] 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
CTC Prefix Search[10] 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
TABLE IV: CER and Noun Recall of GPT-4o on AISHELL-2 Streaming ASR Error Correction. The ASR model is a conformer-based AED model trained on AISHELL-2 and decoded with the Attention Rescore Decoding Method using a chunk size of 16.
LLM Correction Error Metrics Performance Metrics
sub del insert Change CER Noun Recall
5.26 0.17 0.14 - 5.57 77.64
GPT-4o 4.47 0.29 0.19 -0.62 (11%) 4.95 79.66
@@ IV-B The Dataset Configuration
We select three datasets as evaluation benchmarks, including the Chinese datasets AISHELL-1 [23] and AISHELL-2 [24], and the English dataset LibriSpeech [25], aiming to evaluate performance on both Chinese and English data. The LibriSpeech test set is partitioned into two subsets: test-clean, containing high-quality audio, and test-other, comprising noisier audio samples.
IV-C The Model Configuration
@@
In our experiment, we adopt a conformer-based attention-encoder-decoder (AED) ASR model trained with the WeNet toolkit [28], following the U2++[29] method, across three datasets. For AISHELL-1 and LibriSpeech, a non-streaming model architecture is employed, and four decoding strategies are evaluated: Attention, Attention Rescoring, CTC Greedy Search, and CTC Prefix Search.
To further assess the performance of streaming models, experiments are conducted on AISHELL-2 using the Attention Rescoring decoding method with a chunk size of 16. The Conformer encoder comprises 12 blocks with an attention dimension of 256, 4 attention heads, and 2048 linear units, incorporating relative positional encoding and Swish activation. The decoder is implemented as a bidirectional Transformer with three forward and three backward layers. The network follows a hybrid CTC/attention architecture, with the CTC loss weight set to 0.3.
@@
For LLMs, we select GPT-4o [30] and DeepSeek-V2 [31], both demonstrating strong multilingual capabilities in Chinese and English. The decoding temperature is set to 0.2, and the top-$p$ parameter is set to 0.8.
@@
TABLE V: Framework Analysis Results on AISHELL-1 with DeepSeek v2 and Attention Decoding Method; The ASR model is a conformer-based AED model trained on AISHELL-1
LLM Correction Error Metrics Performance Metrics Cost Tokens (in/out)
sub del insert CER Change Noun Recall